# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 5.1.3 Estimation and aggregating across time in event-studies
$Y_{i,t}=\theta_{t}+\eta_{i}+\sum_{k=1}^{g-2}\beta_{k}\left(\mathbf{1}\{G_{i}=g\}\cdot\mathbf{1}\{t=k\}\right)+\sum_{k=g}^{T}\beta_{k}\left(\mathbf{1}\{G_{i}=g\}\cdot\mathbf{1}\{t=k\}\right)+\varepsilon_{i,t}$ (5.3)
@@
This regression produces identical estimates to those obtained “by hand” via (5.2): $\widehat{\beta}_{t}=\widehat{ATT}(t)$. It also generates (point-wise) confidence intervals based on clustered standard errors as discussed in Section 3.3. An additional issue in event study inference, however, involves the fact that we are now estimating many treatment effect parameters, often to study the evolution of $ATT(t)$ over time. Thus, when we compare across event study estimates, we are conducting many hypothesis tests, and the usual normal critical values used to construct confidence intervals do not account for these. Asymptotically correct inferences about the entire event study curve requires “inflating” critical values to perform a multiple hypothesis test adjustment. In Figure 3, the thick black bars represent the standard pointwise confidence intervals from clustered standard errors at the county level, while the red line shows uniform confidence bands that cover the 95% confidence interval for the entire treatment path of the event study coefficients after accounting for multiple testing. These are produced by default in the Callaway2021 statistical packages, using a multiplier bootstrap procedure to compute critical values of the $\sup$-$t$ test statistic. Alternatively, one can construct these using the estimated variance-covariate matrix of all $\widehat{\beta}_{t}$’s paired with the Montiel-Olea_Plagborg-Moller-2018-JAE simulation procedure. Alternative bootstrap procedures, such as the nonparametric bootstrap, the multiplier bootstrap, and the weighted/Bayesian bootstrap, can also be used to compute the $\sup$-$t$ critical values that account for multiple testing. [^19] The $\sup$-$t$ critical value governs the width of the uniform confidence band that yields simultaneous coverage probabilities for a given confidence level (Montiel-Olea_Plagborg-Moller-2018-JAE). In our context, its main idea is constructing asymptotically valid critical values for the entire event-study trajectory based on the maximum of all t-statics (one for each event-time considered). This procedure avoids the conservativeness of other traditional multiple-testing corrections, such as Bonferroni. See Montiel-Olea_Plagborg-Moller-2018-JAE for a general discussion.
@@
A final estimation issue arises when targeting aggregations of the $ATT(t)$s. For example, the average treatment effect in the post-period, $ATT_{\text{avg}}=\frac{1}{T-(g-1)}\sum_{t=g}^{T}ATT(t)$, is a convenient scalar measure that improves statistical precision, especially when $ATT(t)$ is relatively constant. The easiest way to get an estimate $\widehat{ATT}_{\text{avg}}$ is just to construct it from the $2\times 2$ $\widehat{ATT}(t)$ building block estimates. Standard post-estimation commands achieve this if the event study estimates come from a regression, and newer DiD packages report this parameter automatically. A common shortcut, however, is to run a second regression that replaces the event study dummies with the treatment status dummy. $D_{i,t}=\mathbf{1}\{D_{i}=1\}\times\mathbf{1}\{t\geq g\}$:
$Y_{i,t}=\theta_{t}+\eta_{i}+\beta^{OLS}D_{i,t}+\varepsilon_{i,t}$ (5.4)
@@
Unfortunately, the (weighted) least squares estimator $\widehat{\beta}^{OLS}$ does not generally equal the $\widehat{ATT}_{\text{avg}}$. The reason is that $\widehat{\beta}^{OLS}$ is equivalent to first collapsing the multiple-periods data to averages in the post- and pre-periods and then estimating a $2\times 2$ DiD on the resulting means. This, in turn, is the same as subtracting the average pre-period $\tau_{-k}$ estimate (including the zero at time $g-1$) from the average post-period $\widehat{ATT}(t)$ estimate. This implies that when we want to interpret both $\beta_{t}$s from (5.3) and $\beta^{OLS}$ from (5.4), we would need to strengthen Assumption PT-ES to hold in every time period, and not only in the post-treatment periods. [^20] The decision to estimate each $\widehat{ATT}(t)$ relative to period $t=g-1$ comes directly from the choice to define PT that way. When one is comfortable in assuming parallel trends in every period, one can typically form more efficient estimators than those discussed above. See Marcus_Santanna_2021 for a discussion To the extent that the gap in mean outcomes over the whole pre-period differs from the gap in outcomes in period $t=g-1$, the two summary parameters will not be equal. In Figure 3, the two summary parameters are quite different. $\widehat{ATT}_{\text{avg}}$ equals -0.78 while $\widehat{\beta}^{OLS}$ equals -2.60.212121Interestingly, this quantity is almost identical to our estimate of $ATT(2014)$, but rather than representing anything reassuring, it comes from the offsetting effects of positive pre-period estimates and small post-period ones.
@@ 5.1.4 Covariates in event studies
Another advantage of seeing event studies as collections of $2\times 2$ building blocks is that all the tools for incorporating covariates from Section 4 immediately apply to each event study estimate. In fact, the only difference is that instead of using “short-differences” $\Delta Y_{i,t=2}$, one would now use “long-differences”, $Y_{i,t}-Y_{i,t=g-1}$. This would imply that using the regression-adjustment procedure would require estimating a working model for $\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|D_{i}=0,X_{i}]$ for each time $t$. The propensity score working model used to construct the IPW DiD estimate (4.6), on the other hand, is exactly the same as in a $2\times 2$ analysis of the same groups. Since the DR DiD estimation procedure builds on both RA and IPW procedures, it would involve estimating different outcome-regression working models for each time $t$. We also note that the potential pitfalls of controlling for covariates in a TWFE specification still apply with multiple periods and actually become more complex (Caetano2024).
For completeness and ease of access, we list the RA, IPW, and DR estimands for $ATT(t)$,
@@
Notes: This figure shows the corresponding event study estimates with the inclusion of covariates. The outcome variable is the crude mortality rate for adults ages 20-64, and the covariates include the percentage of the county population that is female, the percentage of the county population that is white, the percentage of the county population that is Hispanic, the unemployment rate, the poverty rate, and county-level median income. The point estimate is reported by the circles, and both point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines.
@@ 5.2 Staggered treatment adoption ($\mathbf{G_{\#}\times T}$)
When treatment start dates can vary across units, we need to allow the potential outcomes, and thus the target parameters and identifying assumptions, to reflect this richer notion of treatment. We therefore index potential outcomes by the time treatment begins, $g$: $Y_{i,t}(g)$; and use $Y_{i,t}(\infty)$ to denote never-treated potential outcomes. [^22] Let $\mathbf{0}_{s}$ and $\mathbf{1}_{s}$ be $s$-dimensional vectors of zeros and ones, respectively, and denote the potential outcome for unit $i$ at time $t$ if first exposed to the treatment at time $g$ by $Y_{i,t}(\mathbf{0}_{g-1},\mathbf{1}_{T-g+1})$, and denote by $Y_{i,t}(\mathbf{0}_{T})$ the outcome if untreated by time $t=T$. We discussed the two-period treatment this way in section 3.1 when we defined potential outcomes as a function of the period one and period two treatment. These treatment paths define the potential outcomes we work with: $Y_{i,t}(g)=Y_{i,t}(\mathbf{0}_{g-1},\mathbf{1}_{T-g+1})$ and $Y_{i,t}(\infty)=Y_{i,t}(\mathbf{0}_{T})$. Writing potential outcomes as functions of treatment paths helps with transparency regarding causal parameters of interest and the DiD design we are in. When treatment can turn on and off, writing potential outcomes in terms of the entire path becomes crucial to avoid “hidden” assumptions that rule out treatment effect heterogeneity and dynamics. Due to space constraints, we do not cover these cases in this article. $G_{i}$ denotes each unit’s treatment date, and with some abuse of terminology, we call units not exposed to treatment by period $T$ the “never-treated” group. [^23] In practice, “never-treated” really means “not observed to be treated by $t=T$.” Given more data, units untreated at $T$ could, in many cases, take up the treatment. In fact, this is the case with the Medicaid expansion analysis in which we use data through 2019 but include states that expanded Medicaid in 2020, 2021, and 2023 as “never treated” alongside states that have not expanded as of 2024. Finally, we use $\mathcal{G}$ to represent the set of all treatment times (rows of Table 1 in our example). With these modifications, we can map potential outcomes to observed outcomes using a generalization of (3.1):
@@
Like before, Assumption NA-S imposes that treatment effects are zero in all pre-treatment periods as a consequence of units not acting on the potential knowledge of future treatment dates before they are actually exposed to treatment. We maintain this assumption throughout this section.
Finally, we assume that a “never-treated” group always exists in our staggered DiD setup. If all units are eventually treated, we drop all the data from when the last cohort is treated, so the last-treated cohort becomes the “never-treated” cohort, and $T$ here denotes the number of available periods in the subset of the data that we will use in our analysis. This is essentially without loss of generality because, under standard DiD assumptions, we cannot identify any $ATT$ for periods where all units are treated. We also dropped data from units treated in the first available period, $G_{i}=1$, as such treatment group does not have any pre-treatment data, preventing us from conducting a DiD analysis.
Figure 5 plots average weighted mortality data by the year of Medicaid expansion ($G_{i}$) and time. As in Table 2 and Figure 2, these are all the means necessary to calculate a staggered DiD estimate. Like in Section 5.1, we only report population-weighted results for brevity throughout this section.
@@
Notes: This figure shows county population-weighted average mortality rates for adults ages 20-64 by expansion year from 2009 to 2019.
@@ 5.2.1 Building block parameters with staggered adoption
Staggered treatment timing affects the structure of a DiD analysis because it changes the definition of treatment. Until now, we have used counties in the 2014 expansion states as the only treatment group and have represented them with “the” treatment dummy, $D_{i}$. But Table 1 shows that as of 2019, there were four different groups of expansion states defined by whether they expanded Medicaid in 2014, 2015, 2016, or 2019. Therefore, $D_{i}$ is not rich enough to capture the relevant definition of treatment groups in staggered setups because there are many treatment groups, not just two. Fortunately, we can use the treatment timing notation, $G_{i}$, to define $ATT$ parameters, parallel trends assumptions, and estimators just as we have done so far.
@@ 5.2.2 Identification with staggered designs
Identifying the $ATT(g,t)$s works exactly as in the previous sections because they are just group-specific event studies. Under no anticipation, a set of parallel trends assumptions for $t\geq g$ identifies the causal post-treatment parameters. DiD comparisons for $t<g$ represent differential pre-trends in untreated potential outcomes.
The most important way that staggered DiD changes this approach is that having access to multiple treatment groups with different treatment starting dates allows one to use alternative sets of comparison groups. For example, our Medicaid analysis so far has used counties in states that did not expand Medicaid by 2019 as the comparison group. For estimating, say, the $ATT$ for the 2014 expansion group in 2015, counties in states that did not expand until 2016, 2017, or 2018 could also serve as comparison units. [^24] We can even use data from counties that expanded Medicaid in 2015 as part of our comparison group for $ATT(2014,2015)$, as long as the data for the 2015 cohort that we use comes from 2014 or earlier (which are pre-treatment data for that group). Choosing which comparison groups to use to identify an $ATT(g,t)$ is directly tied to the form of parallel trends. Staggered timing creates many potential parallel trends relative to simple event-studies and especially $2\times 2$ setups. Here, we discuss three types of staggered parallel trends. The first two use either the never-treated units or any not-yet-treated groups as the comparisons for all eventually-treated groups (Callaway2021), and the third option assumes that parallel trends holds in all periods and in all groups, which is something that several recent DiD papers have imposed; see, e.g., deChaisemartin2020; Sun2021; Wooldridge2021; Borusyak2023; Harmon2024.252525 While these three types of parallel trends are fairly intuitive choices, others are possible. For instance, Cengiz2019 use a comparison group of units treated at least $\delta+1$ periods after time $g$. Thus, their parallel trends is tailored to this particular choice. As a result, all of their $ATT(g,t)$ estimates from time $g$ to time $g+\delta$ use the same comparison group. Marcus_Santanna_2021 also discuss other alternative parallel trends assumptions.
@@
Assumption PT-GT-Nev is the analog of the PT assumptions we used in the $2\times T$ design. It uses the never-treated units as the relevant comparison group for all eventually-treated units, and it only imposes parallel trends in post-treatment periods. In our Medicaid application, this would entail using the non-expansion counties as the comparison group for the 2014, 2015, 2016, and 2019 expansion groups. In addition, since Assumption PT-GT-Nev only imposes parallel trends for the future, the farthest we can go into pre-treatment periods is $t=g-1$, which will serve as the only (justifiable) baseline period. More formally, under Assumption PT-GT-Nev, it is straightforward to show that, for post-treatment periods
@@
which shows that $ATT(g,t)$ is identified (Callaway2021; Sun2021). Note that (5.5) highlights that we are essentially back to a $2\times 2$ design when it comes to learning about $ATT(g,t)$: it only leverages data from two periods, $t$ (post) and $g-1$ (pre), and two treatment groups, $G_{i}=g$ (treated) and $G_{i}=\infty$ (comparison).
Assumption PT-GT-NYT allows one to use not only the never-treated units but any group of units that are not-yet-treated by time $t$. In our Medicaid example, we could now use non-expansion counties and 2016 and 2019 expansion counties as comparison groups when estimating $ATT(2014,2015)$. Based on Assumption PT-GT-NYT, Callaway2021 has shown that we can identify the $ATT(g,t)$ for post-treatment periods $t\geq g$ by262626deChaisemartin2020 have derived the same results but restricting attention to instantaneous treatment effects, that is, $ATT(g,g)$’s. They do allow for treatment turning on and off, but also impose that being exposed to a treatment today does not affect outcomes tomorrow (a no-carryover assumption). See deChaisemartin2023b-intertemporal-treatments for some extensions.
@@
Once again, this result highlights identifying the $ATT(g,t)$’s in post-treatment periods is a matter of going back to $2\times 2$ setups: it only leverages data from two periods, $t$ (post) and $g-1$ (pre), and two treatment groups, $G_{i}=g$ (treated), and $G_{i}>t$ (comparison). [^27] We note that more general results are also possible. In fact, for any not-yet-treated group $g^{\prime}>t$, it is easy to show that, under Assumption PT-GT-NYT, for $t\geq g$, $ATT(g,t)=\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|G_{i}=g]-\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|G_{i}=g^{\prime}].$ One can then flexibly combine the different comparison groups by leveraging user-specified or efficiency-oriented weights. Chen_SantAnna_Xie_effientDiD_2024 discuss how to efficiently explore all the information implied by the identification assumptions to form semiparametrically efficient DiD estimators, i.e., estimators that asymptotically enjoy the shortest possible (theoretically-justified) confidence intervals without making strong functional form or model-based assumptions related to error terms such as homoskedasticity and restrictions on serial dependence.
Finally, Assumption PT-GT-all allows the use of any not-yet-treated units as a comparison group, as well as any pre-treatment period as a baseline period. For instance, to identify the ATT for 2015 expansion counties, we can now use the never-treated, 2016 and 2019 expansion groups, and use any or all of the years from 2009 to 2014 as a baseline. Based on Assumption PT-GT-all, it is easy to show that, for any pre-treatment period $t_{\text{pre}}<g$ and any not-yet-treated group $g^{\prime}>t$, we can identify the $ATT(g,t)$ for group $g$’s post-treatment periods $t\geq g$ by
$ATT(g,t)$ $=\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t_{\text{pre}}}|G_{i}=g]-\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t_{\text{pre}}}|G_{i}=g^{\prime}].$ (5.7)
Again, even this more complex estimand (5.7) maps back to the $2\times 2$ DiD setup, as it leverages two periods, $t$ (post) and $t_{\text{pre}}$ (pre), and two groups, $G_{i}=g$ (treated) and $G_{i}=g^{\prime}$ (comparison). This representation makes it clear that, in practice, we can leverage various pre-treatment periods and not-yet-treated comparison groups to characterize the $ATT(g,t)$ under Assumption PT-GT-all. We can also combine several of these to form an $ATT(g,t)$ estimand that uses more data Wooldridge2021; Gardner2021; Liu2022; Borusyak2023; Chen_SantAnna_Xie_effientDiD_2024. An intuitive estimand that naturally extends (5.6) and allows us to use more pre-treatment data is given by

### v2

@@ 5.1.3 Estimation and aggregating across time in event-studies
$Y_{i,t}=\theta_{t}+\eta_{i}+\sum_{k=1}^{g-2}\beta_{k}\left(\mathbf{1}\{G_{i}=g\}\cdot\mathbf{1}\{t=k\}\right)+\sum_{k=g}^{T}\beta_{k}\left(\mathbf{1}\{G_{i}=g\}\cdot\mathbf{1}\{t=k\}\right)+\varepsilon_{i,t}.$ (5.3)
@@
This regression produces identical estimates to those obtained “by hand” via (5.2): $\widehat{\beta}_{t}=\widehat{ATT}(t)$. It also generates (point-wise) confidence intervals based on clustered standard errors, as discussed in Section 3.3. An additional issue in event study inference, however, involves the fact that we are now estimating many treatment effect parameters. Thus, when we compare across event study estimates, we are conducting many hypothesis tests, and the usual normal critical values used to construct confidence intervals do not account for these. Asymptotically correct inferences about the entire event study curve require “inflating” critical values to perform a multiple hypothesis test adjustment. In Figure 3, the thick black bars represent the standard pointwise confidence intervals from clustered standard errors at the county level, while the red line shows uniform confidence bands that cover the 95% confidence interval for the entire treatment path of the event study coefficients after accounting for multiple testing. These are produced by default in the Callaway2021 statistical packages, using a multiplier bootstrap procedure to compute critical values of the $\sup$-$t$ test statistic. Alternatively, one can construct these using the estimated variance-covariate matrix of all $\widehat{\beta}_{t}$’s paired with the Montiel-Olea_Plagborg-Moller-2018-JAE simulation procedure. Alternative bootstrap procedures, such as the nonparametric bootstrap, the multiplier bootstrap, and the weighted/Bayesian bootstrap, can also be used to compute the $\sup$-$t$ critical values that account for multiple testing. [^19] The $\sup$-$t$ critical value governs the width of the uniform confidence band that yields simultaneous coverage probabilities for a given confidence level (Montiel-Olea_Plagborg-Moller-2018-JAE). In our context, its main idea is constructing asymptotically valid critical values for the entire event-study trajectory based on the maximum of all t-statics (one for each event-time considered). This procedure avoids the conservativeness of multiple-testing corrections such as Bonferroni’s. See Montiel-Olea_Plagborg-Moller-2018-JAE for a general discussion.
@@
A final estimation issue arises when targeting aggregations of the $ATT(t)$’s. For example, the average treatment effect in the post-period, $ATT_{\text{avg}}=\frac{1}{T-(g-1)}\sum_{t=g}^{T}ATT(t)$, is a convenient scalar measure that improves statistical precision, especially when $ATT(t)$ is relatively constant. The easiest way to get an estimate of $\widehat{ATT}_{\text{avg}}$ is just to construct it from the $2\times 2$ $\widehat{ATT}(t)$ building block estimates. Standard post-estimation commands achieve this if the event study estimates come from a regression, and newer DiD packages report this parameter automatically. A common shortcut, however, is to run a second regression that replaces the event study dummies with the treatment status dummy. $D_{i,t}=\mathbf{1}\{D_{i}=1\}\times\mathbf{1}\{t\geq g\}$:
$Y_{i,t}=\theta_{t}+\eta_{i}+\beta^{OLS}D_{i,t}+\varepsilon_{i,t}.$ (5.4)
@@
Unfortunately, the (weighted) least squares estimator $\widehat{\beta}^{OLS}$ does not generally equal the $\widehat{ATT}_{\text{avg}}$. The reason is that $\widehat{\beta}^{OLS}$ is equivalent to first collapsing the multiple-periods data to averages in the post- and pre-periods and then estimating a $2\times 2$ DiD on the resulting means. This, in turn, is the same as subtracting the average pre-period $\tau_{-k}$ estimate (including the zero at time $g-1$) from the average post-period $\widehat{ATT}(t)$ estimate. This implies that interpreting both $\beta_{t}$s from (5.3) and $\beta^{OLS}$ from (5.4) requires Assumption PT-ES to hold in every time period, not just in the post-treatment periods. [^20] The decision to estimate each $\widehat{ATT}(t)$ relative to period $t=g-1$ comes directly from the choice to define PT that way. If parallel trends holds in every period, one can typically form more efficient estimators than those discussed above. See Marcus_Santanna_2021 for a discussion To the extent that the gap in mean outcomes over the whole pre-period differs from the gap in outcomes in period $t=g-1$, the two summary parameters will not be equal. In Figure 3, the two summary parameters are quite different: $\widehat{ATT}_{\text{avg}}$ equals -0.70 while $\widehat{\beta}^{OLS}$ equals -2.53.212121Interestingly, this quantity is almost identical to our estimate of $ATT(2014)$, but rather than representing anything reassuring, it comes from the offsetting effects of positive pre-period estimates and small post-period ones.
@@ 5.1.4 Covariates in event studies
Another advantage of seeing event studies as collections of $2\times 2$ building blocks is that all the tools for incorporating covariates from Section 4 immediately apply to each event study estimate. In fact, the only difference is that instead of using “short-differences,” $\Delta Y_{i,t=2}$, one would now use “long-differences,” $Y_{i,t}-Y_{i,t=g-1}$. This would imply that using the regression-adjustment procedure would require estimating a working model for $\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|D_{i}=0,X_{i}]$ for each time $t$. The propensity score working model used to construct the IPW DiD estimate (4.6), on the other hand, is exactly the same as in a $2\times 2$ analysis of the same groups. Since the DR DiD estimation procedure builds on both RA and IPW procedures, it would involve estimating different outcome-regression working models for each time $t$. We also note that the potential pitfalls of controlling for covariates in a TWFE specification still apply with multiple periods and actually become more complex (Caetano2024).
For completeness and ease of access, we list the RA, IPW, and DR estimands for $ATT(t)$:
@@
Notes: This figure shows event study estimates that include covariates. The outcome variable is the crude mortality rate for adults ages 20-64, and the covariates include 2013 values of the percentage of the county population that is female, the percentage of the county population that is white, the percentage of the county population that is Hispanic, the unemployment rate, the poverty rate, and county-level median income. The sample includes 2,200 counties (978 in states that expanded Medicaid by 2014 and 1,222 in states that did not expand Medicaid by 2019). The point estimate is reported by the circles, and both point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines. All procedures use population weights.
@@ 5.2 Staggered treatment adoption ($\mathbf{G_{\#}\times T}$)
When treatment start dates can vary across units, we need to allow the potential outcomes, and thus the target parameters and identifying assumptions, to reflect this richer notion of treatment. We therefore index potential outcomes by the time treatment begins, $g$: $Y_{i,t}(g)$; and use $Y_{i,t}(\infty)$ to denote never-treated potential outcomes. [^22] Let $\mathbf{0}_{s}$ and $\mathbf{1}_{s}$ be $s$-dimensional vectors of zeros and ones, respectively, and denote the potential outcome for unit $i$ at time $t$ if first exposed to the treatment at time $g$ by $Y_{i,t}(\mathbf{0}_{g-1},\mathbf{1}_{T-g+1})$, and denote by $Y_{i,t}(\mathbf{0}_{T})$ the outcome if untreated by time $t=T$. We discussed the two-period treatment this way in section 3.1 when we defined potential outcomes as a function of the period one and period two treatment. These treatment paths define the potential outcomes we work with: $Y_{i,t}(g)=Y_{i,t}(\mathbf{0}_{g-1},\mathbf{1}_{T-g+1})$ and $Y_{i,t}(\infty)=Y_{i,t}(\mathbf{0}_{T})$. Writing potential outcomes as functions of treatment paths helps with transparency regarding causal parameters of interest and the DiD design we are in. When treatment can turn on and off, writing potential outcomes in terms of the entire path becomes crucial to avoid “hidden” assumptions that rule out treatment effect heterogeneity and dynamics. Because of space constraints, we do not cover these cases in this article. We use $G_{i}$ to denote each unit’s treatment date, and with some abuse of terminology, we call units not exposed to treatment by period $T$ the “never-treated” group. [^23] In practice, “never-treated” really means “not observed to be treated by $t=T$.” Given more data, units untreated at $T$ could, in many cases, take up the treatment. In fact, this is the case with the Medicaid expansion. We use data through 2019 but include states that expanded Medicaid in 2020, 2021, and 2023 as “never treated,” alongside states that have not expanded as of 2024. Finally, we use $\mathcal{G}$ to represent the set of all treatment times (rows of Table 1 in our example). With these modifications, we can map potential outcomes to observed outcomes using a generalization of (3.1):
@@
Like Assumption NA, Assumption NA-S imposes that treatment effects are zero in all pre-treatment periods as a consequence of units not acting on the potential knowledge of future treatment dates before they are actually exposed to treatment. We maintain this assumption throughout this section.
Finally, we assume that a “never-treated” group always exists in our staggered DiD setup. If all units are eventually treated, we drop all the data from when the last cohort is treated, so the last-treated cohort becomes the “never-treated” cohort, and $T$ here denotes the number of available periods in the subset of the data that we will use in our analysis. This is essentially without loss of generality, because under standard DiD assumptions, we cannot identify any $ATT$ for periods where all units are treated. We also dropped data from units treated in the first available period, $G_{i}=1$, as such treatment group does not have any pre-treatment data, preventing us from conducting a DiD analysis.
Figure 5 plots average weighted mortality data by the year of Medicaid expansion ($G_{i}$) and time. As in Table 2 and Figure 2, these are all the means necessary to calculate a staggered DiD estimate.
@@
Notes: This figure shows county population-weighted average mortality rates for adults ages 20-64 from 2009 to 2019. There are 978 counties in states in the 2014 expansion group, 171 counties in the 2015 expansion group, 93 counties in the 2016 expansion group, 140 counties in the 2019 expansion group, and 1,222 counties that did not expand Medicaid by 2019.
@@ 5.2.1 Building block parameters with staggered adoption
Staggered treatment timing affects the structure of a DiD analysis because it changes the definition of treatment. Until now, we have used counties in the 2014 expansion states as the only treatment group represented with a single treatment dummy, $D_{i}$. But Table 1 shows that as of 2019, there were four different groups of expansion states defined by whether they expanded Medicaid in 2014, 2015, 2016, or 2019. Therefore, $D_{i}$ is not rich enough to capture the relevant definition of treatment groups in staggered setups, because there are many treatment groups, not just two. Fortunately, we can use the treatment timing notation, $G_{i}$, to define $ATT$ parameters, parallel trends assumptions, and estimators, just as we have done so far.
@@ 5.2.2 Identification with staggered designs
Identifying the $ATT(g,t)$’s works exactly as in the previous sections because they are just group-specific event studies. Under no anticipation, a set of parallel trends assumptions for $t\geq g$ identifies the causal post-treatment parameters. DiD comparisons for $t<g$ represent differential pre-trends in untreated potential outcomes.
The most important way that staggered DiD changes this approach is that having access to multiple treatment groups with different treatment starting dates allows one to use alternative sets of comparison groups. For example, our Medicaid analysis so far has used counties in states that did not expand Medicaid by 2019 as the comparison group. For estimating, say, the $ATT$ for the 2014 expansion group in 2015, counties in states that did not expand until 2016, 2017, or 2018 could also serve as comparison units. Choosing which comparison groups to use to identify an $ATT(g,t)$ is directly tied to which form(s) of parallel trends hold. Relative to simple event-studies and especially $2\times 2$ setups, staggered timing creates many potential parallel trends assumptions. Here, we discuss three types of staggered parallel trends. The first two use either the never-treated units or any not-yet-treated groups as the comparisons for all eventually-treated groups (Callaway2021), and the third option assumes that parallel trends holds in all periods and in all groups, which is something that several DiD methods require; see deChaisemartin2020; Sun2021; Wooldridge2021; Borusyak2023; Harmon2024.242424 While these three types of parallel trends are fairly intuitive choices, others are possible. For instance, Cengiz2019 use a comparison group of units treated at least $\delta+1$ periods after time $g$. Thus, their parallel trends is tailored to this particular choice. As a result, all of their $ATT(g,t)$ estimates from time $g$ to time $g+\delta$ use the same comparison group. Marcus_Santanna_2021 also discuss other alternative parallel trends assumptions.
@@
Assumption PT-GT-Nev is the analog of the PT assumptions we used in the $2\times T$ design. It uses the never-treated units as the relevant comparison group for all eventually-treated units, and it imposes parallel trends in post-treatment periods only. In our Medicaid application, this would entail using the non-expansion counties as the comparison group for the 2014, 2015, 2016, and 2019 expansion groups. In addition, since Assumption PT-GT-Nev imposes parallel trends only for the future, the farthest we can go into pre-treatment periods is $t=g-1$, which will serve as the only (justifiable) baseline period. More formally, under Assumption PT-GT-Nev, it is straightforward to show that for post-treatment periods,
@@
which shows that $ATT(g,t)$ is identified (Callaway2021; Sun2021). Note that (5.5) highlights that we are essentially back to a $2\times 2$ design when it comes to learning about $ATT(g,t)$: it leverages data from only two periods, $t$ (post) and $g-1$ (pre), and two treatment groups, $G_{i}=g$ (treated) and $G_{i}=\infty$ (comparison).
Assumption PT-GT-NYT allows one to use not only the never-treated units but any group of units that are not-yet-treated by time $t$. In our Medicaid example, we could now use non-expansion counties and 2016 and 2019 expansion counties as comparison groups when estimating $ATT(2014,2015)$. Using Assumption PT-GT-NYT, Callaway2021 has shown that we can identify the $ATT(g,t)$ for post-treatment periods $t\geq g$ by252525deChaisemartin2020 have derived the same results but restrict attention to instantaneous treatment effects—that is, $ATT(g,g)$’s. They do allow for treatment turning on and off, but also impose that being exposed to a treatment today does not affect outcomes tomorrow (a no-carryover assumption). See deChaisemartin2023b-intertemporal-treatments for some extensions.
@@
Like equation (5.5), this follows a $2\times 2$ set up: it leverages data from only two periods, $t$ (post) and $g-1$ (pre), and two treatment groups, $G_{i}=g$ (treated), and $G_{i}>t$ (comparison). [^26] We note that more general results are also possible. In fact, for any not-yet-treated group $g^{\prime}>t$, it is easy to show that under Assumption PT-GT-NYT, for $t\geq g$, $ATT(g,t)=\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|G_{i}=g]-\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t=g-1}|G_{i}=g^{\prime}].$ One can then flexibly combine the different comparison groups by leveraging user-specified or efficiency-oriented weights. Chen_SantAnna_Xie_effientDiD_2024 discuss how to efficiently explore all the information implied by the identification assumptions to form semiparametrically efficient DiD estimators—that is, estimators that asymptotically enjoy the shortest possible (theoretically justified) confidence intervals without making strong functional form or model-based assumptions related to error terms such as homoskedasticity and restrictions on serial dependence.
Finally, Assumption PT-GT-all allows the use of any not-yet-treated units as a comparison group, as well as any pre-treatment period as a baseline period. For instance, to identify the ATT for 2015 expansion counties, we can now use the never-treated group and the 2016 and 2019 expansion groups, and use any or all of the years from 2009 to 2014 as a baseline. Using Assumption PT-GT-all, it is easy to show that, for any pre-treatment period $t_{\text{pre}}<g$ and any not-yet-treated group $g^{\prime}>t$, we can identify the $ATT(g,t)$ for group $g$’s post-treatment periods $t\geq g$ by
$ATT(g,t)$ $=\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t_{\text{pre}}}|G_{i}=g]-\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}-Y_{i,t_{\text{pre}}}|G_{i}=g^{\prime}],$ (5.7)
which again maps back to the $2\times 2$ DiD setup, as it leverages two periods, $t$ (post) and $t_{\text{pre}}$ (pre), and two groups, $G_{i}=g$ (treated) and $G_{i}=g^{\prime}$ (comparison). This representation makes it clear that, in practice, one can use various pre-treatment periods and not-yet-treated comparison groups to characterize the $ATT(g,t)$ under Assumption PT-GT-all. We can also combine several of these to form an $ATT(g,t)$ estimand that uses more data (Wooldridge2021; Gardner2021; Liu2022; Borusyak2023; Chen_SantAnna_Xie_effientDiD_2024). An intuitive estimand that naturally extends (5.6) and allows us to use more pre-treatment data is given by